Kluth Corporation has two manufacturing departments--Molding and Customizing. The company used the following data at thee beginning of the year to calculate predetermined overhead rates Molding Customizing Total Estimated total machine-hours (MHs) Estimated total fixed manufacturing overhead cost Estimated variable manufacturing overhead cost per MH 1,800 15,800 14,000 $35,000 $ 2.00 $5,040 $40,040 4.00 During the most recent month, the company started and completed two jobs--Job C and Job M. There were no beginning inventories Data concerning those two jobs follow Direct materials Direct labor cost Molding machine-hou Customizing machine-hours Job C Job M $14,700 $ 8,400 $21,600 $ 8,400 2,700 11,300 600 1,200 Requirec Assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in botlh production departments. Further assume that the company uses a markup of 20% on manufacturing cost to establish selling prices Calculate the selling prices for Job C and for Job M. (Do not round intermediate calculations.) Selling price for Job C Selling price for Job M
Answer #1
first we have to compute overhead rate …
Molding Customizing
a Total fixed manufacturing overhead cost = 35000 5040
b Total machine hr. 14000 1800
c=a/b Total fixed overhead per machine hr 2.50 2.80
d Variable manufacturing overhead rate per MH 2 4
e=c+d Total overhead rate = 4.50 6.80
Computation of sales price Job C Job M
A Direct material 14700 8400
B Direct labor 21600 8400
Manufacturing overhead
C Molding =4.5*2700 12150 =4.5*11300 50850
D Customizing =6.8*1200 8160 =6.8*600 4080
E=A+B+C+D Total manufacturing cost 56610 71730
F=E*20% Mark up 20% of cost 11322 14346
G=E+F Sles price 67,932.00               86,076.00
Selling price for Job C     67,932.00
Selling price for Job M     86,076.00